OpenStudy (anonymous):

aDecide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

Can anyone help i dont understand.

OpenStudy (anonymous):

@Preetha @Partycool @texaschic101

OpenStudy (anonymous):

@skullpatrol @Ashleyisakitty @Luigi0210

OpenStudy (anonymous):

b^2 - 4ac is the correct term for factoring. Can you solve the rest?

OpenStudy (anonymous):

I get to 7/25 4

OpenStudy (anonymous):

then i dont know what next

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

Compare your quadratic equation\( 2x^2 + 7x + 3 = 0 \) with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\]

OpenStudy (anonymous):

I get all that part i put in everything right i just get stuck at 7/25 4

hartnn (hartnn):

whats that ? tell me what a,b,c values u got ?

OpenStudy (anonymous):

A=2 B=7 C=3

hartnn (hartnn):

ohh, you mean you got b^2 -4ac = 25 right ? thats it! value of b^2-4ac was asked!

OpenStudy (anonymous):

\[7\sqrt{25}/4\] this is how far i get but the four is on the bottom

hartnn (hartnn):

and since it is a perfect square, the equation can be solved by factoring :)

OpenStudy (anonymous):

ohhh that what iam doing wroung i was using quadric equation thank you again

hartnn (hartnn):

welcome ^_^